# Exploiting Precision Mapping and Component-Specific Feature Enhancement for Breast Cancer Segmentation and Identification

### v4

@@ 5 Conclusion and Future Direction
The primary objective of this research is to detect and segment tumor regions within breast ultrasound images, subsequently categorizing them as benign, malignant, or normal. The objective of this work is to develop an accurate and efficient system for breast cancer tumor segmentation and classification, aiming to improve diagnosis and treatment outcomes for patients. The proposed segmentation model utilizes an InceptionResNet-based LinkNet framework with an intelligent dual-attention mechanism to precisely segment the tumor region. Leveraging spatial and self-attention mechanisms across multiple layers, the DCNNIMAF classification framework enables accurate classification of breast cancer types or the absence of cancerous conditions. The proposed models have excelled in performance, in comparison to existing works. In segmentation tasks, they showcase exceptional accuracy, IoU score, and Dice coefficient score. Furthermore, the classification metrics reveal impressive accuracy, precision, F1-score, and recall rates. Future work could extend the framework’s utility to other medical imaging modalities, facilitating the detection and classification of abnormalities beyond breast ultrasound images.

### v5

@@ 5 Conclusion and Future Direction
In this paper, we propose two novel DL frameworks to overcome the two important hurdles in the diagnosis of breast cancer using ultrasound imaging: PMAD-LinkNet for segmentation and CSFEC-Net for classification. The segmentation framework leverages the Precision Mapping Mechanism (PMM) for locations to be spatially mapped automatically, thus allowing accurate tumor boundary refinement in the presence of morphological variations and disparate image quality. The classification framework features the Component-Specific Feature Enhancement Module (CSFEM), which utilizes multi-level attention to emphasize discriminative features between benign, malignant, and normal tissues.
While the suggested frameworks have demonstrated remarkable results, it does not stop the late work possible in this area. The generalizability of the models could be improved by collecting larger samples with greater diversity and by augmenting data multimodally. For instance, incorporate images of mammograms and obtain magnetic resonance imaging (MRI) data. Second, using explainable AI techniques could go a long way toward improving the interpretability of model decisions and hence increasing trust and clinical adoption. Third, real-time implementations and optimizations should be researched and developed for deployment in edge devices so that it could enable more wide accessibility in non-resource settings. Finally, adapting the frameworks to all other possible imaging modalities and applications in medicine would further show their versatility and impact on healthcare. In doing so, we will make those avenues possible that will enhance the applicability and effectiveness of proposed methods toward improving early detection and treatment outcomes concerning breast cancer and many more critical diseases.